OpenStudy (anonymous):

A number a is a root of P(x) if and only if the remainder, when dividing the polynomial by x - a, equals zero. A. True B. False

OpenStudy (mathmale):

What do YOU think? What are your reasons for your choice?

OpenStudy (anonymous):

its true

OpenStudy (anonymous):

@mathmale

OpenStudy (mathmale):

What are your reasons for your answer choice? Perhaps you could answer that by giving an example of long division or synthetic div. where the remainder is zero.

OpenStudy (michele_laino):

Hint: we can apply this theorem about polynomilas: "GIven the polynomial \(P(x)\) and the binomial \(x-a\), then we can find two other polynomials \(q(x)\) and \(r(x)\), such that, the subsequent condition holds:" \[P\left( x \right) = q\left( x \right)\left( {x - a} \right) + r\left( x \right)\] where degree of q(x) is less than degree of P(x)

OpenStudy (michele_laino):

oops... about polynomials*

OpenStudy (mathmale):

Yes (Michele). Referring to that Theorem, what conclusion can we draw if r(x)=0?